\begin{thebibliography}{10}

%\cite{Kuzenko:2011ti}
\bibitem{KT} 
  S.~M.~Kuzenko and S.~J.~Tyler,
  ``Complex linear superfield as a model for Goldstino,''
  JHEP {\bf 1104}, 057 (2011)
  [arXiv:1102.3042 [hep-th]].
  %%CITATION = ARXIV:1102.3042;%%

\bibitem{VA}
D.~V.~Volkov and V.~P.~Akulov,
``Possible universal neutrino interaction,''
 % \href{http://www.jetpletters.ac.ru/ps/1766/article_26864.shtml}
  {JETP Lett.\  {\bf 16}, 438 (1972)}   
  [Pisma Zh.\ Eksp.\ Teor.\ Fiz.\   {\bf 16}, 621 (1972)]; 
  ``Is the neutrino a Goldstone particle?,''
 % \href{http://dx.doi.org/10.1016/0370-2693(73)90490-5}%
  {Phys.\ Lett.\  B {\bf 46}, 109 (1973)}.
 %%CITATION = ZFPRA,16,621;%%

%\cite{Ivanov1977}
%\bibitem{Ivanov1977}
\bibitem{Ivanov}
E.~A.~Ivanov and A.~A.~Kapustnikov,
 ``Relation between linear and nonlinear realizations of supersymmetry,''
 %\href{http://www.slac.stanford.edu/spires/find/hep/www?r=JINR-E2-10765}
 {%
 Preprint JINR-E2-10765, June 1977};
 %%CITATION = JINR-E2-10765;%%
%\bibitem{Ivanov1978}
%E.~Ivanov and A.~Kapustnikov, 
``{General relationship between linear and
  nonlinear realisations of supersymmetry},''
  %\href{http://dx.doi.org/10.1088/0305-4470/11/12/005}
  {
  {J.\ Phys.\  A} {\bfseries 11},  2375 (1978)};
%\bibitem{Ivanov1982}
%E.~Ivanov and A.~Kapustnikov, 
``{The nonlinear realisation structure of models
  with spontaneously broken supersymmetry},''
%  \href{http://dx.doi.org/10.1088/0305-4616/8/2/004}%
  {{J.\ Phys.\ G}  {\bfseries 8}, 167 (1982)}.


\bibitem{Uematsu:1981rj}
T.~Uematsu and C.~K.~Zachos,
  ``{Structure of phenomenological Lagrangians for broken supersymmetry},''
 % \href{http://dx.doi.org/10.1016/0550-3213(82)90431-X}
  {Nucl.\ Phys.\  B {\bf 201}, 250 (1982).}
  %%CITATION = NUPHA,B201,250;%%



\bibitem{SamuelWess1983}
S.~Samuel and J.~Wess, ``{A superfield formulation of the non-linear
  realization of supersymmetry and its coupling to supergravity},''
  %\href{http://dx.doi.org/10.1016/0550-3213(83)90622-3}
  {{Nucl.\ Phys.\ B}
  {\bfseries 221,} 153 (1983)}.


%\cite{Farakos:2015vba}
\bibitem{FHKvU} 
  F.~Farakos, O.~Hulik, P.~Ko{c}i and R.~von Unge,
  ``Non-minimal scalar multiplets, supersymmetry breaking and dualities,''
  arXiv:1507.01885 [hep-th].
  %%CITATION = ARXIV:1507.01885;%%

%\cite{Rocek:1978nb}
\bibitem{Rocek} 
  M.~Ro\v{c}ek,
  ``Linearizing the Volkov-Akulov model,''
  Phys.\ Rev.\ Lett.\  {\bf 41}, 451 (1978).
  %%CITATION = PRLTA,41,451;%%
  %117 citations counted in INSPIRE as of 15 juil. 2015

%\cite{Kuzenko:2011tj}
\bibitem{KT0}
  S.~M.~Kuzenko and S.~J.~Tyler,
  ``On the Goldstino actions and their symmetries,''
  JHEP {\bf 1105} (2011) 055
  [arXiv:1102.3043 [hep-th]].
  %%CITATION = ARXIV:1102.3043;%%
  %12 citations counted in INSPIRE as of 15 Jul 2015

%\cite{Butter:2011vg}
\bibitem{BK} 
  D.~Butter and S.~M.~Kuzenko,
  ``A dual formulation of supergravity-matter theories,''
  Nucl.\ Phys.\ B {\bf 854}, 1 (2012)
  [arXiv:1106.3038 [hep-th]].
  %%CITATION = ARXIV:1106.3038;%%

\end{thebibliography}

\RequirePackage{ifpdf} 
\ifpdf
\documentclass[pdftex,12pt,a4paper]{article}  
\else
\documentclass[12pt,a4paper]{article} 
\fi

\parskip=\medskipamount

\usepackage[mathscr]{euscript} 	
\usepackage{bbm} 
\usepackage{amsfonts,amsmath,amssymb}
\ifpdf
\usepackage[pdftex]{graphicx}	
\usepackage[usenames,pdftex]{color}		
\usepackage[pdftex,bookmarks]{hyperref}
\hypersetup{bookmarksopen,bookmarksnumbered,
		pdfnewwindow=true,
		pdfauthor = {Simon J. Tyler},
		pdftitle = {Comments on the complex linear Goldstino superfield},
		pdfsubject = {},
		pdfkeywords = {Supersymmetry, Goldstino},
		pdfcreator = {LaTeX with hyperref package},
		pdfproducer = {pdflatex}
}
\else
\usepackage[dvips]{graphicx}
\usepackage[usenames,dvips]{color}
\usepackage[dvips,bookmarks,colorlinks=false]{hyperref}
\fi
\def\a{\alpha}

\def\G{\Gamma}

\def\k{\kappa}

\def\q{\theta}

\def\S{\Sigma} 

\def\X{\Xi}
\newcommand{\eps}{\varepsilon}

\newcommand{\da}{{\dot{\alpha}}}
\newcommand{\db}{{\dot{\beta}}}

\def\ds1{\ensuremath{\mathbbm{1}}}
\newcommand{\rmd}{{\rm d}}

\newcommand{\rmi}{{\rm i}}

\newcommand{\cN}{{\cal N}}

\newcommand{\pd}{\partial}

\def\intx{\int\!\!{\rmd}^4x\,}

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\bea}{\begin{eqnarray}}
\newcommand{\eea}{\end{eqnarray}}

\newcommand{\Db}{{\bar{D}}}

\begin{document}                        
\begin{titlepage}

\begin{flushright} July, 2015 \end{flushright}

\vspace{5mm}

\begin{center}
{\large \bf  Comments on the complex linear Goldstino superfield}
\end{center}
\begin{center}
{\large  
{Sergei M.\ Kuzenko}
and 
{Simon J.\ Tyler}

\vspace{5mm}

\footnotesize{
{\it School of Physics M013, The University of Western Australia\\
35 Stirling Highway, Crawley W.A. 6009, Australia}}  

\vspace{2mm}}
\end{center}

\vspace{5mm}

\pdfbookmark[1]{Abstract}{abstract_bookmark}
\begin{abstract}
\baselineskip=14pt
It is shown that the complex linear Goldstino superfield 
recently proposed in  arXiv:1507.01885 is obtained from the one originally constructed in 
arXiv:1102.3042 by a field redefinition. 
\end{abstract}
\vfill

\end{titlepage}


Four years ago, we constructed 
the Goldstino model  \cite{KT} described by 
a {\it modified} complex linear superfield $\S$,
\begin{align} \label{mCL:constraint}
	-\frac14 \Db^2 \S = f\,, \qquad f = {\rm const}\,.
\end{align}
Here $f$ is a parameter of mass dimension 2 which, 
without loss of generality, can be chosen to be real.
To describe the Goldstino dynamics, $\S$ was   subject to 
the following nonlinear constraints:
\begin{align}
	\S^2 &= 0~, \label{1st constraint} \\
	-\frac{1}{4} \S\Db^2D_\a\S &= f D_\a\S~. \label{2nd constraint}
\end{align}
The constraint \eqref{1st constraint} tells us that $\S$ is nilpotent.
The form of the Goldstino action coincides with the free action for the complex linear superfield,
\bea
	S[\S,\bar\S] 
	= - \intx\!\rmd^2\q\rmd^2\bar\q\, \S\bar\S  ~.
	\label{Gaction}
\eea

It was also shown in \cite{KT} that all known Goldstino superfields can be obtained as composite constructed from spinor covariant derivatives of
$\S$ and its conjugate $\bar \S$.\footnote{%
This property and the universality  \cite{VA,Ivanov,Uematsu:1981rj}
of the Goldstino \cite{VA} implies that any model for supersymmetry breaking 
can be described in terms of $\S$ and its conjugate.} 
Such constructions make use of
the spinor superfields $\X_\a$
and its conjugate ${\bar \X}_\da$ defined by 
\begin{align} \label{SW_as_Derivative}
\X_\a = \frac{1}{\sqrt2}D_\a \bar \S\,, \qquad
	\bar\X_\da = \frac{1}{\sqrt2}\Db_\da\S\ .
\end{align}
Making use of the constraints \eqref{mCL:constraint}, \eqref{1st constraint} 
and \eqref{2nd constraint}, we can readily uncover 
those constraints which are obeyed by the above spinor superfields. 
They have the form (with  $2\k^2=f^{-2}$)
\begin{subequations}\label{SW_Constraint}
\begin{align}
\label{SW_Constraint1}
	{\bar D}_\da {\bar \X}_\db 
		&= \k^{-1} \eps_{\da \db}\,, \\
\label{SW_Constraint2}
	D_\a {\bar \X}_\da
		&= 2 \rmi \k {\bar \X}^\db \pd_{\a \db} {\bar \X}_\da  
\end{align}
\end{subequations}
and are exactly the constraints given in 
\cite{SamuelWess1983}, so we recognise $\X_\a$ as the Samuel-Wess superfield.
In particular, for the Goldstino superfield $\S$ 
the solution to the constraints \eqref{mCL:constraint}, \eqref{1st constraint} 
and \eqref{2nd constraint} in terms of $\bar \X_\da$ 
is very simple:
\begin{align} \label{mCL-SW-soln}
	2f\S = \bar{\X}^2 \ ,
\end{align}
and the action \eqref{Gaction} takes the form 
\bea
	S[\S,\bar\S] 
	= - \frac{1}{4f^2}\intx\!\rmd^2\q\rmd^2\bar\q\, \X^2 \bar \X^2  ~.
	\label{Gaction2}
\eea

One option for a composite Goldstino superfield that was not discussed in \cite{KT} 
was recently explored in \cite{FHKvU}. It is defined in terms of the Samuel-Wess superfield as
\begin{align}
\Gamma = -\frac{\k^2}{\sqrt{8}} \bar D_\da ( \bar \X^\da \X^2)~.
\end{align}
Making use of \eqref{SW_Constraint} gives
\begin{align}
\G = \frac{\k}{\sqrt{2}} \X^2 \Big\{ 1-\k^2 {\rm i}
\bar \X_\da \partial^{\da \a}\X_\a \Big\}~.
\label{10}
\end{align}
By construction, $\G$ is   an {\it ordinary} complex linear superfield,
\begin{align}
\bar D^2 \G =0\,,
\end{align}
and it follows from \eqref{Gaction2} and \eqref{10} that it has an identical action to $\S$,
\begin{align}
S[\S,\bar\S] = S[\G , \bar \G]~,
\end{align}
where $S[\G , \bar \G]$ is obtained from \eqref{Gaction} 
by replacing $\S \to \G$.
As a consequence of \eqref{10}, $\G$ it is nilpotent,
\begin{align}
\G^2 =0~.
\end{align}
Equations \eqref{SW_Constraint1} and \eqref{SW_Constraint2}
imply that $\G$ satisfies a nonlinear constraint  
that was not discussed  in \cite{FHKvU}
\begin{align}
-\frac14\G\bar D^2\bar\G = f\G ~.
\label{CL-2nd-constraint}
\end{align}
This is structurally identical to the second constraint of Rocek \cite{Rocek},
which is natural since $\G$ is on-shell equivalent to Rocek's constrained chiral superfield \cite{FHKvU}.
As this constraint mixes $\G$ with its conjugate, it leads to complicated solutions for its components.
This is in contradistinction to the constraint \eqref{2nd constraint} that only depends on $\S$.
The solutions to the constraint \eqref{CL-2nd-constraint} and all of the calculations in this paragraph 
can be found in the ancilliary Mathematica notebook attached to \cite{KT0}.

It is interesting to note that in the \emph{modified} complex linear superfield of \cite{KT} 
the constraints mean that the Goldstino arises from the normally physical fermion 
$G_\a \propto \left.D_\a \bar\S\right|$. 
In the complex linear superfield $\G$, the constraints mean that the Goldstino arises 
from a normally auxiliary spinor $G_\a \propto \left.D_\a \G\right|$. 
This was discussed in \cite{FHKvU} and is now easily seen from \eqref{10} 
when rewritten using \eqref{SW_as_Derivative} as 
\begin{align}
\G = \bar\S - \frac{\k}{\sqrt8}(\Db_\da\S)(\Db^\da\bar\S) 	~ .
\end{align}

In conclusion, we would like to point out that the $\S$-realisation \cite{KT}
and the $\G$-realisation \cite{FHKvU} for complex linear
Goldstino superfield are to some extent complementary. 
The former is characterised by the holomorphic-like  constraint  \eqref{2nd constraint} 
that is very useful for applications. The latter has been deduced from
models allowing spontaneous supersymmetry breaking.  
It is worth mentioning that the constraint \eqref{mCL:constraint}
is the only way to describe $\cN=1$ anti-de Sitter supergravity
using a non-minimal scalar multiplet \cite{BK}.


\begin{footnotesize}
\providecommand{\href}[2]{#2}
\begingroup\raggedright
\endgroup
\end{footnotesize}

\end{document}